If 7.8 grams of hydrogen reacts with 13.7 grams of oxygen to produce water, how many grams of water would be produced according to the following equation? Round your answer to two decimal places.



2H2 + O2 --> 2H2O
1
Expert's answer
2019-03-08T05:01:31-0500

n(H2)= 7.8g"/ 2 {g \over mol} = 3.9 mol"

n(O2)= "13.7g/(15.99 {g \over mol}*2) = 0.428 mol"

2H2 + O2 "\to" 2H2O

n(H2O)= 2*n(O2)= 0.428*2=0.856 mol

m(H2O)=n(H2O)*M(H2O) = "0.856mol*18.01 {g \over mol} = 15.42 g"